# Safety-Critical Control with Guaranteed Lipschitz Continuity via Filtered Control Barrier Functions

### v3

@@ II Definitions and Preliminaries
Lemma 1 ([7]).
Let $b:\mathbb{R}^{n}\to\mathbb{R}$ be a continuously differentiable function and let $\bm{x}(t)$ be a trajectory of system (1) over $t\in[t_{0},t_{1}]$. If the time function $b(\bm{x}(t))$ satisfies $\dot{b}(\bm{x}(t))\geq-\alpha(b(\bm{x}(t)))$ for all $t\in[t_{0},t_{1}]$, where $\alpha$ is a class $\mathcal{K}$ function of its argument, and $b(\bm{x}(t_{0}))\geq 0$, then $b(\bm{x}(t))\geq 0$ for all $t\in[t_{0},t_{1}]$.
@@ III Problem Formulation and Approach
A control policy is feasible if (9) and (2) are strictly satisfied $\forall t\in[0,T].$ In this paper, we consider the following problem:
@@ IV-B Filtered Control Barrier Functions (FCBF)
where $\bm{\pi}(t)\coloneqq[\pi_{1}(t),\dots,\pi_{m_{a}q}(t)]^{{}^{\top}}\in\mathbb{R}^{m_{a}q}$ denotes an auxiliary state with $\pi_{j}(t)\in\mathbb{R},j\in\{1,...,m_{a}q\}$. $\nu\in\mathbb{R}^{q}$ is an auxiliary input for (13), $F:\mathbb{R}^{m_{a}q}\to\mathbb{R}^{m_{a}q}$ and $G:\mathbb{R}^{m_{a}q\times q}\to\mathbb{R}^{m_{a}q}$ are locally Lipschitz. For simplicity, we just build up the connection between auxiliary variables and the system as $\bm{\pi}_{1:q}=\bm{u}_{f},\bm{\pi}_{(q+1):2q}=\bm{u}^{(1)}_{f},\dots,\bm{\pi}_{((m_{a}-1)q+1):m_{a}q}=\bm{u}^{(m_{a}-1)}_{f},\bm{u}^{(m_{a})}_{f}=\bm{\nu}$. $\bm{\pi}_{i:j}$ refers to the entries of vector $\bm{\pi}$ from index $i$ to $j$ and $\bm{u}_{f}^{(i)}$ represents the $i^{\text{th}}$ derivative of $\bm{u}_{f}$ with respect to time $t$. A more general construction of system (13) requires that the auxiliary state $\bm{u}_{f}$ has relative degree $m_{a}$ with respect to system (13). In order to guarantee safety, we first define $\psi_{0,f}(\bm{x},\bm{\pi})\coloneqq\psi_{m}(\bm{x},\bm{u}_{f})$ based on (6), where the relative degree of $\psi_{0,f}(\bm{x},\bm{\pi})$ with respect to system (13) is $m_{a}$. We then define a sequence of functions $\psi_{i,f}:\mathbb{R}^{n}\to\mathbb{R},\ i\in\{1,...,m_{a}\}$ as
@@ IV-C Optimal Control with FCBFs
where $\left\|\cdot\right\|$ denotes the 2-norm of a vector, $D(\cdot)$ is a strictly increasing function of its argument and $T>0$ denotes the ending time. $p\left\|\bm{x}(T)-\bm{x}_{e}\right\|^{2}$ denotes state convergence similar to (8). To minimize $\left\|\bm{u}_{f}\right\|$, we can perform input–output linearization for (13) and use the CLF introduced in  [28, Eq. 24] to minimize $\bm{u}_{f}^{\top}\bm{u}_{f}$. We can then formulate the CLFs (in Def. 6 and  [28, Eq. (24)]), HOCBFs (in Thm. 3 and Cor. 1) and FCBFs (in Def. 7 and Cor. 1) as constraints of the QP with cost function (23) to realize safety-critical control.
@@ V Case Study and Simulations
In this section, we consider the unicycle model with the dynamics given by (24) for Problem 1, which is more realistic than the simplified unicycle model used in Sec. IV-A and in the case study of [20]. We use MATLAB’s quadprog solver to solve the QP at each time step, and integrate the system dynamics using the ode45 function. All computations are performed on a computer equipped with an Intel® Core™ i7-11750F CPU @ 2.50 GHz. The average computation time for each QP is less than 0.01 s.
@@
$\underbrace{\begin{bmatrix}\dot{x}(t)\\ \dot{y}(t)\\ \dot{\theta}(t)\\ \dot{v}(t)\\ \dot{u}_{f1}(t)\\ \dot{u}_{f2}(t)\end{bmatrix}}_{(\dot{\bm{x}}(t),\dot{\bm{u}}_{f}(t))}=\begin{bmatrix}v(t)\cos{(\theta(t))}\\ v(t)\sin{(\theta(t))}\\ u_{f1}(t)\\ \frac{u_{f2}(t)}{M}\\ -\frac{1}{\tau}u_{f1}(t)\\ -\frac{1}{\tau}u_{f2}(t)\end{bmatrix}+\begin{bmatrix}0&0\\ 0&0\\ 0&0\\ 0&0\\ \frac{1}{\tau}&0\\ 0&\frac{1}{\tau}\end{bmatrix}\underbrace{\begin{bmatrix}\nu_{1}(t)\\ \nu_{2}(t)\end{bmatrix}}_{\bm{\nu}(t)}.$ (25)
@@
$\begin{split}&\psi_{i,0}^{\text{min}}=u_{fi}-u_{i,\text{min}},\psi_{i,0}^{\text{max}}=u_{i,\text{max}}-u_{fi}\\ &\psi_{i,1}^{\text{min}}(u_{fi},\nu_{i})\coloneqq\dot{\psi}_{i,0}^{\text{min}}(u_{fi},\nu_{i})+k_{i,1}^{\text{min}}\psi_{i,0}^{\text{min}},\\ &\psi_{i,1}^{\text{max}}(u_{fi},\nu_{i})\coloneqq\dot{\psi}_{i,0}^{\text{max}}(u_{fi},\nu_{i})+k_{i,1}^{\text{max}}\psi_{i,0}^{\text{max}},\end{split}$ (29)
where $\alpha_{i,1}^{\text{min}}(\cdot),\alpha_{i,1}^{\text{max}}(\cdot)$ are defined as linear functions and $i\in\{1,2\}$. To satisfy the state convergence, we define a CLF $V(\bm{x}(t),\bm{\pi}(t))=(10(\theta(t)-\theta_{d})+u_{f1}(t)+u_{f2}(t))^{2}$ with $\theta_{d}=atan2(\frac{y_{d}-y(t)}{x_{d}-x(t)}),c_{1}=c_{2}=1$ to stabilize $\theta(t)$ to $\theta_{d}$ and $u_{f1}(t)+u_{f2}(t)$ to 0 and formulate the relaxed constraint in (7) as
@@
Other parameters are $c_{3}=10,M=1650kg,u_{f1}(0)=u_{f2}(0)=0,Q=10^{5},k_{1}=k_{2}=10,k_{1,1}^{\text{min}}=k_{2,1}^{\text{min}}=k_{1,1}^{\text{max}}=k_{2,1}^{\text{max}}=\alpha$. The above method is referred to as FCBF. For the first benchmark, we only include constraints from HOCBFs (27), input (26) and CLFs (30) with $V(\bm{x}(t))=(\theta(t)-\theta_{d})^{2}$ and refer to this as HOCBF method (note that the decision variables in the cost function (31) are $\bm{u}(t),\delta(t)$). For the second benchmark, we build upon the first benchmark by adding $0.1(u_{1}(t)-u_{1}(t-\Delta t))^{2}+0.1(u_{2}(t)-u_{2}(t-\Delta t))^{2}$ to the cost function. This term encourages gradual changes by penalizing the difference between the control inputs at consecutive time steps. We refer to the second benchmark as sp-HOCBF (smoothness-penalized HOCBF).

### v4

@@ II Definitions and Preliminaries
Lemma 1 (Comparison Lemma, e.g., [11]).
Let $y:\mathbb{R}^{n}\to\mathbb{R}$ be a continuously differentiable function and let $\bm{x}(t)$ be a trajectory of system (1) over $t\in[t_{0},t_{1}]$. If the time function $y(\bm{x}(t))$ satisfies $\dot{y}(\bm{x}(t))\geq-\alpha(y(\bm{x}(t)))$ for all $t\in[t_{0},t_{1}]$, where $\alpha$ is a class $\kappa$ function of its argument, and $y(\bm{x}(t_{0}))\geq 0$, then $y(\bm{x}(t))\geq 0$ for all $t\in[t_{0},t_{1}]$.
@@ III Problem Formulation and Approach
A control policy is feasible if (9) and (2) are satisfied $\forall t\in[0,T].$ In this paper, we consider the following problem:
@@ IV-B Filtered Control Barrier Functions (FCBF)
where $\bm{\pi}(t)\coloneqq[\pi_{1}(t),\dots,\pi_{m_{a}q}(t)]^{{}^{\top}}\in\mathbb{R}^{m_{a}q}$ denotes an auxiliary state with $\pi_{j}(t)\in\mathbb{R},j\in\{1,...,m_{a}q\}$. $\nu\in\mathbb{R}^{q}$ is an auxiliary input for (13), $F:\mathbb{R}^{m_{a}q}\to\mathbb{R}^{m_{a}q}$ and $G:\mathbb{R}^{m_{a}q}\to\mathbb{R}^{m_{a}q\times q}$ are locally Lipschitz. For simplicity, we just build up the connection between auxiliary variables and the system as $\bm{\pi}_{1:q}=\bm{u}_{f},\bm{\pi}_{(q+1):2q}=\bm{u}^{(1)}_{f},\dots,\bm{\pi}_{((m_{a}-1)q+1):m_{a}q}=\bm{u}^{(m_{a}-1)}_{f},\bm{u}^{(m_{a})}_{f}=\bm{\nu}$. $\bm{\pi}_{i:j}$ refers to the entries of vector $\bm{\pi}$ from index $i$ to $j$ and $\bm{u}_{f}^{(i)}$ represents the $i^{\text{th}}$ derivative of $\bm{u}_{f}$ with respect to time $t$. A more general construction of system (13) requires that the auxiliary state $\bm{u}_{f}$ has relative degree $m_{a}$ with respect to system (13). In order to guarantee safety, we first define $\psi_{0,f}(\bm{x},\bm{\pi})\coloneqq\psi_{m}(\bm{x},\bm{u}_{f})$ based on (6), where the relative degree of $\psi_{0,f}(\bm{x},\bm{\pi})$ with respect to system (13) is $m_{a}$. We then define a sequence of functions $\psi_{i,f}:\mathbb{R}^{n}\to\mathbb{R},\ i\in\{1,...,m_{a}\}$ as
@@ IV-C Optimal Control with FCBFs
where $\left\|\cdot\right\|$ denotes the 2-norm of a vector, $D(\cdot)$ is a strictly increasing function of its argument and $T>0$ denotes the ending time. $p\left\|\bm{x}(T)-\bm{x}_{e}\right\|^{2}$ denotes state convergence similar to (8). To minimize $\left\|\bm{u}_{f}\right\|$, we can perform input–output linearization for (13) and use the CLF introduced in  [27, Eq. 24] to minimize $\bm{u}_{f}^{\top}\bm{u}_{f}$. We can then formulate the CLFs (in Def. 6 and  [27, Eq. (24)]), HOCBFs (in Thm. 3 and Cor. 1) and FCBFs (in Def. 7 and Cor. 1) as constraints of the QP with cost function (23) to realize safety-critical control.
@@ V Case Study and Simulations
In this section, we consider the unicycle model with the dynamics given by (24) for Problem 1, which is more realistic than the simplified unicycle model used in the case study of [26]. We use MATLAB’s quadprog solver to solve the QP at each time step, and integrate the system dynamics using the ode45 function. All computations are performed on a computer equipped with an Intel® Core™ i7-11750F CPU @ 2.50 GHz. The average computation time for each QP is less than 0.01 s.
@@
$\underbrace{\begin{bmatrix}\dot{x}(t)\\ \dot{y}(t)\\ \dot{\theta}(t)\\ \dot{v}(t)\\ \dot{u}_{f_{1}}(t)\\ \dot{u}_{f_{2}}(t)\end{bmatrix}}_{(\dot{\bm{x}}(t),\dot{\bm{u}}_{f}(t))}=\begin{bmatrix}v(t)\cos{(\theta(t))}\\ v(t)\sin{(\theta(t))}\\ u_{f_{1}}(t)\\ \frac{u_{f_{2}}(t)}{M}\\ -\frac{1}{\tau}u_{f_{1}}(t)\\ -\frac{1}{\tau}u_{f_{2}}(t)\end{bmatrix}+\begin{bmatrix}0&0\\ 0&0\\ 0&0\\ 0&0\\ \frac{1}{\tau}&0\\ 0&\frac{1}{\tau}\end{bmatrix}\underbrace{\begin{bmatrix}\nu_{1}(t)\\ \nu_{2}(t)\end{bmatrix}}_{\bm{\nu}(t)}.$ (25)
@@
$\begin{split}&\psi_{i,0}^{\text{min}}=u_{f_{i}}-u_{i,\text{min}},\psi_{i,0}^{\text{max}}=u_{i,\text{max}}-u_{f_{i}}\\ &\psi_{i,1}^{\text{min}}(u_{f_{i}},\nu_{i})\coloneqq\dot{\psi}_{i,0}^{\text{min}}(u_{f_{i}},\nu_{i})+k_{i,1}^{\text{min}}\psi_{i,0}^{\text{min}},\\ &\psi_{i,1}^{\text{max}}(u_{f_{i}},\nu_{i})\coloneqq\dot{\psi}_{i,0}^{\text{max}}(u_{f_{i}},\nu_{i})+k_{i,1}^{\text{max}}\psi_{i,0}^{\text{max}},\end{split}$ (29)
where $\alpha_{i,1}^{\text{min}}(\cdot),\alpha_{i,1}^{\text{max}}(\cdot)$ are defined as linear functions and $i\in\{1,2\}$. To satisfy the state convergence, we define a CLF $V(\bm{x}(t),\bm{\pi}(t))=(10(\theta(t)-\theta_{d})+u_{f_{1}}(t)+u_{f_{2}}(t))^{2}$ with $\theta_{d}=atan2(\frac{y_{d}-y(t)}{x_{d}-x(t)}),c_{1}=c_{2}=1$ to stabilize $\theta(t)$ to $\theta_{d}$ and $u_{f_{1}}(t)+u_{f_{2}}(t)$ to 0 and formulate the relaxed constraint in (7) as
@@
Other parameters are $c_{3}=10,M=1650kg,u_{f_{1}}(0)=u_{f_{2}}(0)=0,Q=10^{5},k_{1}=k_{2}=10,k_{1,1}^{\text{min}}=k_{2,1}^{\text{min}}=k_{1,1}^{\text{max}}=k_{2,1}^{\text{max}}=\alpha$. The above method is referred to as FCBF. For the first benchmark, we only include constraints from HOCBFs (27), input (26) and CLFs (30) with $V(\bm{x}(t))=(\theta(t)-\theta_{d})^{2}$ and refer to this as HOCBF method (note that the decision variables in the cost function (31) are $\bm{u}(t),\delta(t)$). For the second benchmark, we build upon the first benchmark by adding $0.1(u_{1}(t)-u_{1}(t-\Delta t))^{2}+0.1(u_{2}(t)-u_{2}(t-\Delta t))^{2}$ to the cost function. This term encourages gradual changes by penalizing the difference between the control inputs at consecutive time steps. We refer to the second benchmark as sp-HOCBF (smoothness-penalized HOCBF).